OpenStudy (thuru7):

If the Density of a Mixture of Helium and Oxygen gasses in 25 degrees Celcious temperature , 8.314x10^5 Pa , is 0.835 g/dm^-3 , Find the Percentage of mass of Helium included in the mixture? [ R.A.M {He=4 , O=16}

OpenStudy (aaronq):

use: PV=nRT, sub in \(n=\dfrac{m}M{}\) and \(d=\dfrac{m}{V}\) and isolate for M, which is the average of the molar masses of the gases. (1) \(M_{ave}=\dfrac{dRT}{P}\) The \(M_{ave}\) is made up of the molar mass and % abundance of each gas. (2) \(M_{ave}=M_{He}*\%He+M_{O_2}*\%O_2\) Their percentages equal 100%, so 1 (3) \(\%He+\%O_2=1\) plug numbers given in the question into (1), \(M_{ave}=29.944~g/mol\) combine (2) and (3), plug in molar masses, solve for %He \(M_{ave}=M_{He}*\%He+M_{O_2}*(1-\%He)=29.944~g/mol\) \(\%He=0.0734=\color{red}{7.34\%}\) which is the mass percent of He of the mixture.